Minutes of the 31st SOHO SWT Meeting

University of Manchester, Manchester, UK 9 August 2000

2

Contents

1 Welcoming Address 5 2 Action Items 5 3 Agree Agenda and Action Revision 5 4 Spacecraft Status 6 5 S/C Offpoint Request 6 6 MEDOC Campaign/Station Keeping Manoeuvre 6 7 TRACE 6 8 5 Years of SOHO Celebrations 7 9 Workshops and Meetings 7 10 SOHO Archive Status 7 11 New JOP with ACE, EUVE, Ulysses 7 12 AOB 7 Annex 1: Agenda 9 Annex 2: List of Participants 11 Annex 3: Spacecraft Status 13 Annex 4: S/C Offpoint Request 21 Annex 5: MEDOC Campaign/Station Keeping Manoeuvre 37 Annex 6: Resolution from the SWT in support of TRACE 39 Annex 7: 5 Years of SOHO Celebrations 41 Annex 8: Workshops and Meetings 43 Annex 9: Java Interface to the SOHO Archive 47 Annex 10: JOP Proposal with ACE, EUVE, Ulysses 51 4

1 Welcoming Address

B. Fleck welcomed the attendees. All instruments except ERNE were represented — J. Torsti had sent his apologies for not being able to attend.

2 Action Items

Action 31-1: On J. Gurman: Details on offpointing manoeuvre to be distributed (see annex 4) and a proposal to be sent to the Program Office. Feedback from risk analysis to be distributed.

Action 31-2: On B. Fleck: Place and time of 5-year celebrations to be selected.

Action 31-3: On PIs: Suggest material for “5-year booklet”.

3 Agree Agenda and Action Revision

No changes to the agenda (see annex 1). List of attendees - annex 2.

Actions Revision

Action 30-1: Closed^{1} On PIs: Provide L. Sánchez Duarte^{2} with default data policy for archival data, if any restrictions are desired.

No inputs received – no restrictions applied.

Action 30-2: Closed^{1} On PIs: Provide updated information for the SOHO Bibliography & Publications database^{3}, with particular emphasis on Ph.D. Theses.

The TRACE bibliography will be added.

CDS, CELIAS, ERNE, GOLF, SUMER, UVCS, MDI in good shape.

For COSTEP, LASCO, EIT the NASA ADS will be queried, however all JGR and GRL papers will be missed. Searches for SWAN and VIRGO are impossible.

Action 30-3: Closed^{1} On PIs: Request team members to send paper copies of papers in journals (other than ApJ and A&A) to the SOHO PS office^{4}.

Received one package from SUMER and one from a UVCS team member.

1These items are always ongoing but will not be kept as open actions.

2lsanchez@esa.nascom.nasa.gov

3Submit information to soho@esa.nascom.nasa.gov

4SOHO PS Office, NASA GSFC, Mail Code 682.3, SOHO EAF, Bldg 26 Room G-1, Greenbelt, MD 20771

4 Spacecraft Status

B. Fleck presented the spacecraft status (annex 3) on behalf of M. Chaloupy, J.-P. Olive, and H. Schweitzer.

The only main event since SWT-30 was the 14 July 2000 X 5.7 flare, resulting in a 1.75% degradation of the solar arrays. The output is nevertheless well within the budget, and total consumption is less than 2/3 of the power provided.

The star swapping features of the new gyroless software performed flawlessly during the proton storm (several star swaps and recoveries were experienced).

A brief overview of the offpointing manoeuvre (“profile”) capabilities was given.

5 S/C Offpoint Request

J. Gurman presented the case for an offpointing manoeuvre to enable the calculation of a new flatfield for EIT (see annex 4).

Summary of discussion: The maximum pointing excursion will be about 40 arcminutes, performed in Normal Mode, with Medium Rate telemetry. Golf/MDI are willing to accept the data loss. MDI would benefit from the manoeuvre if high rate telemetry is possible. The offpointing may be useful to CELIAS/SEM. Space weather conditions will be taken into account before the profile is performed. The likely time frame will be end of November/after the MEDOC campaign. There will be a possibility of saying no at a later stage.

The manoeuvre is to be proposed to the Program Office, separate from other manoeuvres/station keeping activities. The risk analysis is to be distributed to the PIs for final approval (action 31-1).

6 MEDOC Campaign/Station Keeping Manoeuvre

B. Fleck presented the situation concerning the next MEDOC campaign and the need for a station keeping/momentum management manoeuvre (see annex 5).

It was agreed to propose a new date (27 October) for the manoeuvre, keeping the MEDOC campaign in week 44/45.

7 TRACE

A. Title informed about the funding problems for TRACE operations, which will impact SOHO/TRACE collaborations.

Since first light, TRACE has participated in more SOHO campaigns than any SOHO instrument!

The SWT agreed that a resolution in support of TRACE should be issued, to be drafted by J. Gurman (see annex 6).

8 5 Years of SOHO Celebrations

B. Fleck initiated a discussion about the various options available to mark the the SOHO 5th anniversary (see annex 7).

It was agreed that a combination of “celebratory talks”, public lectures and science news should be arranged, with invitations to ESA, NASA and industry representatives. Place and time is yet TBD (action 31-2).

Arranging a workshop along the lines of “From Solar Min to Solar Max – Half a Solar Cycle with SOHO” was regarded as too much work considering the time available, and left for later. Volunteers are welcome!

All PIs should suggest material for a “5-year booklet” (action 31-3).

9 Workshops and Meetings

A list of known meetings can be found in annex 8. See the “Meetings” page^{5} on the SOHO Web server for updates.

10 SOHO Archive Status

B. Fleck presented the status of the SOHO archive, and announced the availability of the JAVA interface through e.g. the SOHO home page (see also annex 9). The need for “analysis cookbooks” to complement the data archive was emphasized. The guaranteed duration of the archive is 10 years after the end of the mission.

11 New JOP with ACE, EUVE, Ulysses

A new JOP proposed by E. Möbius (JOP leader) was presented by R. Lallement (annex 10). The local contact at the EOF would be J. Raymond. The JOP was approved.

12 AOB

SWT-32 will be held Thursday 14 December 2000 at Stanford, in conjunction with the AGU Fall Meeting, 15-19 December 2000.

5http://sohowww.nascom.nasa.gov/meetings/

8

Annex 1: Agenda

9

Annex 2: List of Participants

11

Attendees SOHO SWT-31

Name  Institute/Experiment  Address
Brekke, P.  ESA/SSD  NASA/GSFC, Code 682.3 (pbrekke@esa.nascom.nasa.gov)
Curdt, W.  MPAE/SUMER  (curdt@linmpi.mpg.de)
Dere, K.  NRL/LASCO  (kdere@solar.stanford.edu)
Fleck, B.  ESA/SSD  NASA/GSFC, Code 682.3 (bfleck@esa.nascom.nasa.gov)
Fröhlich, C.  PMOD-WRC/VIRGO  CH-7260 Davos-Dorf
Gabriel, A.  IAS/GOLF  Bat 121, Univ. Paris XI, F-91405 Orsay Cedex (gabriel@ias.fr)
Galvin, A.B.  Univ. New Hampshire/CELIAS SSC, Morse Hall, Durham
Gurman, J. NASA/GSFC/EIT (gurman@gsfc.nasa.gov)
Harrison, R.A. RAL/CDS  RAL, Didcot, Oxon, OX11 OQX, UK
Haugan, S.V.H. ESA/SSD  NASA/GSFC Code 682.3 (shaugan@esa.nascom.nasa.gov)
Kunow, H.  Univ. Kiel/COSTEP  Inst. für Exp. u. Angewandte Physik, Univ. Kiel,
Lallement, R.  CNRS/SWAN  BP 3, Verrieres-Le-Buisson, F-91371
Noci, G.  Univ. Firenze/UVCS  Largo Fermi 5, 50125, Firenze, Italy (noci@arcetri.astro.it)
Smith, P.  CfA/UVCS  (plsmith@solar.stanford.edu)
Title, A.  LPARL/TRACE  LPARL, 3251 Hanoverstr., Palo Alto, CA 94304, USA
Turck-Chieze  Saclay/GOLF  Service d’Astrophysique/DAPNIA, CE Saclay,

Annex 3: Spacecraft Status

13 20

Annex 4: S/C Offpoint Request

21

Annex 5: MEDOC Campaign/Station Keeping Manoeuvre

37

Annex 6: Resolution from the SWT in support of TRACE

39

Annex 7: 5 Years of SOHO Celebrations

41

Annex 8: Workshops and Meetings

43 46

Annex 9: Java Interface to the SOHO Archive

47

Annex 10: JOP Proposal with ACE, EUVE, Ulysses

51